### v1

# Squeezing Quantum States in Three-Dimensional Twisted Crystals –ِ

Squeezing Quantum States in Three-Dimensional Twisted Crystals
A fundamental idea in wave mechanics is that propagation in a periodic medium can be described by Bloch waves whose conserved crystal momenta define their transformations when displaced by the set of discrete lattice translations. In ordered materials where incommensurate spatial periods compete, this general principle is rendered ineffective, often with dramatic consequences. Examples are crystals with broken symmetries from charge or spin density waves [monceau_electronic_2012, takayama_continuum_1980], quasiperiodic lattices that produce diffraction patterns with crystallographically forbidden point symmetries [yu_dodecagonal_2019, rotenberg_electronic_2004], and stacks of two-dimensional lattices with a relative rotation (twist) between layers [andrei_marvels_2021]. In special cases when there is a small difference between the competing periods, a useful work-around is a continuum description where a periodic long-wavelength field produces Bragg scattering that coherently mixes short-wavelength carrier waves [brazovskii_exact_1980, bistritzer_moire_2011]. In this work, we advocate an alternative approach to study three-dimensional twisted crystals that replaces their spectrally congested momentum-space Bloch band structures [cea_twists_2019, wu_three-dimensional_2020] with a representation using squeezed coherent states in a Fock space of free-particle vortex states [bliokh_theory_2017]. This reorganization of the Hilbert space highlights the crucial role of the Coriolis force in the equations of motion that leads to unconventional phase space dynamics and edge state structure generic to a family of complex crystals. Screw symmetry of a twisted crystal allows one to separate modes into invariant sectors labelled by a screw eigenvalue [wu_three-dimensional_2020] effectively replacing the plane-wave momentum . However, since the screw operation couples interlayer translation with in-plane rotation, this choice also necessitates abandoning at the outset the conventional Bloch wave representation of the dynamics perpendicular to the screw axis. Instead, each sector inherits a Coriolis coupling to the twist, similar to the 2D motion of a charged particle in a uniform magnetic field, along with an (outward) centrifugal potential. This latter feature distinguishes this system from the related problem in rotating trapped ultracold atomic gases where the centrifugal potential is nearly balanced by a confining potential allowing a faithful mapping to Landau level dynamics from the Coriolis coupling [fletcher_geometric_2021, crepel_geometric_2023, mukherjee_crystallization_2022]. By contrast, in a twisted crystal the centrifugal potential acts to restore nearly “flat” two-dimensional dynamics to a system that would otherwise be spatially confined by the Coriolis deflection. We find that by augmenting the free particle Hamiltonian to include a crystal potential we encounter two classes of quantum states: (i) a spectrally isolated low energy narrow band sector with states that propagate through saddle points in the crystal potential and (ii) a high energy dispersive sector where states ride over the potential landscape and can propagate freely. At lateral boundaries the spectral gap between these two sectors is traversed by helical edge states in which reversed orbital angular momentum states counterpropagate ballistically in a pattern of winding edge channels. In the coherent state representation presented below all these features can be identified as the generic signatures of wave motion in screw symmetric media.
Fock Space in a Twisted Crystal
In three dimensions, aligning the direction with the screw axis, the free particle kinetic energy can be projected into independent screw sectors with the replacement . Completing the squares on the and components of the momenta, we obtain
Here, contains a -dependent Coriolis coupling to the twist, analogous to the orbital coupling of a charge particle to a uniform magnetic field and a magnetic length . Physically, as a particle moves vertically it encounters a rotating crystal frame which can induce a Coriolis deflection. Since depends on the product , sectors at see reversed effective ’s. For a two-dimensional system in a uniform perpendicular magnetic field the last two terms in Eq. (1) are absent and after separation of variables in mutually commuting left-handed () and right-handed () degrees of freedom the first term reverts to a Hamiltonian for a two-dimensional harmonic oscillator. The momenta have reversed commutation relations and the oscillator Hamiltonian projects entirely into one of the two sectors depending on the field orientation. The oscillator eigenstates form a ladder of discrete states (the Landau levels) that inherit extensive orbital degeneracies from the independent (guiding center) degree of freedom. In rotating systems the situation is more complicated because the centrifugal potential is inevitably present. However, experiments on rotating quantum atomic gases manage to avoid it by balancing using a trapping potential and access the Landau level dynamics [fletcher_geometric_2021, crepel_geometric_2023]. Applications to uniformly twisted crystals introduce three new essential considerations. (i) Both “right” and “left”-handed sectors are accessible in the Hilbert space. A map of the Fock space of number states for the two oscillator quanta 111The mutually commuting operators and where satisfy (see Supplementary Information). is shown in Fig. 1(a). (ii) Propagation along the screw direction disperses the bands in energy, violating the Landau level flat band condition. (iii) The centrifugal potential is present and plays the crucial role of opposing the tendency of the oscillator states to localize around a single (gauge dependent) axis of rotation. We examine these three features in turn. (i) The number states displayed in Fig. 1 are free particle Laguerre-Gauss vortex states [bliokh_theory_2017] which can be written (with ) as
and provide a complete basis where the amplitude for a state with angular momentum and total number is peaked on spiralling paths with mean projected radius . These states are related to superposition states produced in free electron vortex beams [bliokh_theory_2017] and in light beams with orbital angular momentum [allen_orbital_1992]. They are “ultra-localized” around a putative axis of rotation, being confined by the gauge field and have a Gaussian decay in the far field determined by the -dependent magnetic length. (ii) The energy dispersion associated with motion along the stacking direction for a state with definite is given by (completing the square for the last two terms in Eq. (1)) which provides a -dependent effective potential for a lateral two-dimensional projected problem (dashed curves in Fig. 1(b)) and separates the free particle spectrum into overlapping branches with minima at . At these special values, the phases accumulated by the in-plane and out of plane motions exactly cancel giving a state with minimal kinetic energy. (iii) The centrifugal potential has a bilinear form in bosonic raising and lowering operators
which provides a compact expression for the kinetic energy operator in Eq. (1)
The first term is a quadratic form verifying that Eq. (2) recovers the free particle kinetic energy. In the absence of scattering from a twisted potential (whose effects will be examined below) free particles travel in straight lines. Note also that Eq. (2) restores the symmetry in the first term between the ladder operators for the - and - quanta that is necessarily missing from the pure magnetic field problem. This symmetry is broken by the coupling of to in the second term: the Coriolis deflection requires a physical potential that produces a screw axis seen in a sector with through the cross term . Eq. (2) demonstrates the essential role of the centrifugal potential. The kinetic energy operator contains number nonconserving terms that conserve the orbital angular momentum while coupling different number states along decoupled diagonal strings in the Fock space (represented with the bold blue arrows in Fig. 1). Bogoliubov transformations in each string identify the new ground states as superpositions where the total number of quanta is allowed to fluctuate, allowing radial excursions of the amplitudes where the energy penalty for promotion to the higher states is balanced by the energy gained in the centrifugal potential. A representative spectrum of states in the presence of a harmonic confining potential is shown as the bold curves in Fig. 1(b).
State Squeezing at a Saddle Point
In an isotropic medium the level crossings at in Fig. 1(b) are protected by angular momentum conservation. However, a twisted crystal potential lowers this to a discrete rotational symmetry that preserves the screw symmetry. In a typical case the relevant potential has the lateral period of an emergent moir superlattice [andrei_marvels_2021, bistritzer_moire_2011] but the physical consequences are captured by an even simpler situation with state mixing at an isolated saddle point where the principal axes of the saddle precess with the pitch : which has twofold rotational symmetry but preserves continuous screw symmetry. mixes angular momentum branches and within a common screw sector producing a pattern of avoided crossings. A useful limiting case is for the extremal states where the coherent states have the minimum allowed radius for each value of the angular momentum and the saddle potential gaps out the free particle spectrum (Fig. 1(c)) [crepel_microscopic_2020]. Explicitly, the action of on the number states is which has the form of a two-mode state-squeezing operator [crepel_geometric_2023]. Because of the quadrupolar form, the scattered states exchange two units of angular momentum with the twisted lattice and therefore the angular momentum is conserved mod 2. This isolates two sets of low energy bands in which the superposed angular momentum states are solely even- or odd- integer valued. The saddle point potential also acts to destablize the lowest frequency modes in each angular momentum sector (Bogoliubov branches develop imaginary eigenvalues shown as the red lines in Fig. 1(c) ). This is the signature of unstable trajectories that fall off the saddle point [crepel_geometric_2023]. With open boundary conditions, the radius of the system sets a bound on the maximum accessible angular momentum states . Dispersing branches that saturate this bound occur at the outer boundaries and have no partners at larger radius with which to mix. Consequently these boundary branches are not gapped out and traverse the bulk gaps as shown in Fig. 1(c) and more clearly in Fig. 2(a). Accounting for the independent even- and odd- integer angular momentum families of low energy modes in the bulk, there are two such helical channels at the edge, each containing counterpropagating modes in which opposite angular momenta propagate ballistically in opposite directions along the screw axis. Combinations of these edge modes from the even- and odd- angular momentum families are “barber pole” channels forming pairs of winding trajectories around the boundary. These channels are similar to to the hinge modes at the boundaries of higher order topological insulators [schindler_higher-order_2018, benalcazar_quantized_2017] and indeed the origin in both cases is the quadrupolar order of quantum states in the bulk. Backscattering in these surface channels is strongly suppressed (though in principle it is symmetry allowed for sufficiently large momentum scattering) because of the large mismatch of their screw eigenvalues , analogous to the symmetry protection of surface modes on topological crystalline insulators [fu_topological_2011, kruthoff_topological_2017]. Some features of these quantum states can be understood from classical considerations and are expected to occur in any twisted medium. In Fig. 4(a) we display the solution of the classical equation of motion giving the time evolution of the mean position of a coherent state released from the saddle point with a small initial velocity. We observe trajectories in the form of circulating orbits whose direction is determined by the sign of and which grow in size drifting off the saddle point potential before repeatedly reflecting at turning points on the sidewalls of the saddle potential. The structure of the edge states also follows from a classical model. If we null the kinetic energy, the unstable trajectories away from the saddle point are equilibrated at the sample boundary (Fig. 3(b)). On a cylinder the potential extrema form a pair of classical winding barber pole channels on opposite surfaces. Superpositions of these surface modes correspond to the pairs of edge states found in the even- and odd- angular momenta surface channels in the quantum theory. Removing the restriction allows the radii of the coherent states to adjust in the the saddle potential. Including these processes we again find that the two lowest branches of the coupled spectrum separate from a manifold of higher energy states and generate a pair of helical edge states at the outer surface that traverse this gap (Fig. S4 in the Supplementary Information). Including this relaxation the spectrum contains these fundamental modes along with a pattern of overtones which are the quantized fluctuations of the orbitals around their optimized shapes in the confining potential. These overtones form narrow bands in the bulk each of which are connected to one-sided boundary branches distributed beneath the surface forming a “soft edge” at reduced radii . The lowest energy branches are the fundamental edge states pinned to the exterior boundary found in the “stiff-orbit” model shown in Fig. 2.
Hofstadter Networks
In a three-dimensional twisted crystal each layer sees a potential that is periodic on an emergent moir superlattice scale [andrei_marvels_2021]. In this situation the single saddle point (SP) is replaced by an energy landscape containing a translationally ordered array of compensating SP’s located at sublattice positions repeated on a Bravais superlattice (Fig. 4(a,b)). In the strong field limit, the Coriolis deflection confines states in orbits near the potential extrema. Unstable trajectories that propagate along paths connecting SP’s can escape by propagating along two-dimensional periodic networks [fertig_transmission_1987, chalker_percolation_1988, pal_emergent_2019, de_beule_network_2021]. Importantly, although the crystal potential is invariant under the network lattice translations , the coherent states accumulate path-dependent phases when displaced by ’s making them noncommutative [zak_magnetic_1964, haldane_origin_2018]. Thus the superlattice displacements are generally not identified with the set of magnetic translations [zak_magnetic_1964] that define an effective magnetic unit cell with area which is continuously tunable. The noncommutative structure of the elementary translations provides a local measure for topological order in a band even in the absence of long range crystalline order [bianco_mapping_2011, dereli_bloch_2021]. Here the effective magnetic flux penetrating a single unit cell of the moir superlattice changes as a function of the conserved twist momentum and produces quantum interference in a Hofstadter spectrum whose gap structure depends on the connectivity of the network. Two examples are given in Fig. 4 for models with square and triangular Bravais lattices. In either case the Hofstadter spectra depend on whether the intra-saddle point amplitudes (double well tunneling between a pair of equilibrium states bound near a single SP [novaes_generalized_2003]) or intra-saddle point amplitude [fertig_transmission_1987, chalker_percolation_1988] (paths connecting neighboring SP’s) are dominant. In the latter case the spectrum shows the signature of well developed resonances from weakly coupled intra-cell orbital loops linked by the effective flux. Either limit in Fig. 4 should be physically realizable, being controlled by the relative strengths of the moir crystal field, the kinetic energy penalty for confining states within a moir cell and the strength of the regularizing (pinning) potential. In a typical case, for sufficiently small rotation angle and weak pinning potential, the resonant patterns in Fig. 4(d) and 4(f) are generic. They describe a reconstruction of the low energy degrees of freedom for a nearly free particle in a twisted crystal potential that produces orbital degeneracies at the sublattice nodes of an emergent moir superlattice. Note that the edge state structure for the single SP problem (Fig. 2) reconstructs in the super moir problem so that boundary modes appear within the bulk Hofstadter gaps (Figs. S5 and S6 in the Supplementary Information).
Discussion
The twist-derived gauge coupling in this model is time-reversal symmetric and by itself does not produce a Hall response: opposing contributions to response functions from twist sectors cancel. This can be avoided by biasing the populations of the orbital states with an applied magnetic field, by exciting with circularly polarized light [ji_opto-twistronic_2023] or by spontaneous symmetry breaking from the projected interactions between electrons residing in these narrow bands. Many of the features found here in three dimensions have striking counterparts in twisted two-dimensional few-layer van der Waals materials [andrei_marvels_2021]. Theory suggests and experiments observe bands of low energy degrees of freedom, possibly with topological character, spectrally isolated from remote manifolds of higher energy dispersive states. The conditions under which these low energy bands can develop coherence with the analytic structure of lowest Landau level states has been addressed in recent studies [tarnopolsky_origin_2019, ledwith_vortexability_2023]. A promising approach is to now relate the ideas developed here in three dimensions to the phenomenology of quasi two-dimensional ultrathin materials.
Author Contributions
E.J.M. and V.T.P. conceived the project and contributed to all phases of the analysis. S.A. contributed to the design and interpretion of models for the coherent state spectrum. K.K. carried out analysis of the the Bogoliubov spectra. R.J.S. contributed to symmetry analysis of the model. C.D.B. and M.M.A.Z. analyzed the Hofstadter network. All authors contributed to writing and editing the manuscript.
Acknowledgments
We thank R. Kamien and C. Kane for helpful discussions and V. Crpel and R. Fletcher for useful correspondence. This work was supported by the Department of Energy under grant DE-FG02-84ER45118, Royal Society exchange grant IES/R1/221060 and Singapore Ministry of Education AcRF Tier 2 Grant No. MOE-T2EP50220-0016. V.T.P. is supported by C. Lewandowski’s start-up funds from Florida State University and the National High Magnetic Field Laboratory. The National High Magnetic Field Laboratory is supported by the National Science Foundation through NSF/DMR-2128556 and the State of Florida.
Supplemental Material
Contents
S1 A Tale of Two Reference Frames
S1.1 Coordinate Systems
In a three-dimensional twisted crystal, there are two equivalent frames of reference using which to describe the dynamics. The first is the usual laboratory frame. This is the frame adopted in the main text because it is familiar and intuitive. It takes the perspective of an external observer. In this frame, we can either affix a rectangular coordinate system or a cylindrical coordinate system It seems only sensible to place the origin at the twist axis. The second is the crystal frame, also referred to as the twisted frame, in which the coordinates co-rotate with the crystal. This frame adopts the point of view of the crystal, not the external observer; therefore, it has the advantage that it naturally encodes the helical symmetry of the crystal, and consequently, allows certain crystal potentials to be written in a simplified manner. In this frame, we can also affix a helical coordinate system which is defined relative to the laboratory coordinates as follows [vachagina2014use]:
where is the pitch of the twist (and has dimension of inverse length). This frame is chosen such that a fixed point winds in counterclockwise fashion as increases, as shown in Fig. S1. In the laboratory frame, the Laplacian operator is simply
In the twisted frame, the Laplacian operator is more complicated
This can be derived straightforwardly using repeated applications of the chain rule: we observe that the in-plane component of the Laplacian has the same form in both frames, while the out-of-plane component of the Laplacian is more complicated in the twisted frame, Here, a term related to the angular momentum operator, naturally makes its appearance. The integration measure in both frames is of the same form:
S1.2 Dual Representations of the Free-Particle Hamiltonian
In this work, we work with Schrodinger particles with quadratic dispersion. In any frame, the free-particle Hamiltonian is just where is the mass of the particle. In the laboratory frame, the momentum operator is
Thus, the Hamiltonian in the laboratory frame has the representation
Similarly, the Hamiltonian in the twisted frame has the representation
where we have defined new momentum operators in the twisted frame analogously to those defined in the laboratory frame. It is apparent that commutes with the Hamiltonian, so we can project into a sector of definite eigenvalue of This allows us to replace with a -number: To bring Eq. (S5) into a form similar to Eq. (S6), we define a screw operator in the laboratory frame:
In terms of the screw operator, we have Using the canomical commutation relations we find that
Therefore, Thus, we can also project into subspaces of definite screw eigenvalues of In the end, we have two different representations of the same Hamiltonian with the exact same form:
Despite apparent formal equivalence, the reader is reminded that these two representations are in two completely different frames, and consequently are projected into different physical sectors: in the twisted frame, the Hamiltonian is projected into sectors with plane wavevectors , while in the laboratory frame, the Hamiltonian is projected into sectors with definite screw eigenvalues . Nonetheless, the formal equivalence makes it explicit that both frames are equally useful as long as appropriate care is dedicated when physical interpretation is required. For formal manipulations, only commutation relations matter, and so, both representations give identical results. Consequently, we shall drop the subscripts indicating frames henceforth for brevity whenever there is no risk for confusion (restoring them only when necessary for emphasis), and the reader is free to choose either or to do the projection. In keeping with the main text, we shall choose going forward.
S1.3 Plane Waves versus Screw States
In the twisted frame, we use to label eigenvalues of the operator The associated eigenfunctions are simply plane waves These functions are separable in the sense that they depend only on In the laboratory frame, we use to label eigenvalues of the operator (actually, and are the same operator just expressed in two different frames). In polar coordinates, and this operator is just
It is then straightforward to observe that the eigenfunctions of are of the form
where is an integer labeling the angular momentum. Rewritten in the twisted frame where these eigenvalues have the nice form (notice that this is also an eigenfunction of ). We notice that the screw states differ from the plane waves only in that they encode a helical twist explicitly. Because of that, these states depend on both and in both frames, and consequently, they depend also on angular momentum. In other words, eigenstates of couple the in-plane and out-of-plane degrees of freedom (unless the angular momentum is zero).
S2 Landau-Level Representation of the Free-Particle Hamiltonian
S2.1
In this section, we derive the Landau-level representation of the Hamiltonian. By completing the square, we find
The first two terms remind us of canonical momentum being boosted by a gauge vector field. If we define a magnetic vector potential with an associated magnetic field we find that where is the elementary charge unit. Therefore, we observe that the screw eigenvalue (or the momentum in the -direction) plays the role of a magnetic field that couples to the planar motion in the - plane. For , this magnetic field vanishes because of a decoupling between the out-of-plane and in-plane degrees of freedom. The term is a centrifugal potential that enters the Hamiltonian importantly with a negative sign irrespective of the sign of . The term is an overall shift in energy depending on the screw eigenvalue (or the vertical momentum). Ordinarily, shifts in the spectrum can be disregarded, but we caution against that temptation in this case since this is a relative shift that depends on and not an overall shift of the whole Hamiltonian. The final term measures -projected angular momentum of the wavefunction. For later convenience, we define the cyclotron frequency (this quantity is the same as in the main text, where stands for cyclotron there; here, we instead emphasize the dependence on for later convenience) and magnetic length (this quantity is the same as in the main text; here, we instead emphasize the dependence on for later convenience) in analogy to the Landau-level problem:
We now introduce the left-handed and right-handed operators mentioned in the main text:
The operators are transformed into the operators under exchanging Consequently, these operators degenerate when More seriously, the and operators completely disappear from these expressions. Therefore, the case must always be handled carefully separately. We will return to this at the end of the section. For now, ignoring these operators satisfy the following commutation relations:
Clearly, these four operators are Hermitian. We note that the commutation change sign based on Factors of are omitted in the main text for succinctness, but they need to be considered to obtain the correct algebra. Here, we carefully keep track of all the factors of We are now ready to define ladder operators:
We emphasize that these and operators are defined in a definite sector. So a different gives a different set of operators. If one insists, one can emphasize this dependence by including subscripts: and ; we leave them off to avoid cluttered notion. The factors of have been inserted specifically to ensure the usual ladder commutation relations:
Our task now is to write the Hamiltonian in terms of these ladder operators. To do that, let us rewrite the momentum and position operators in terms of the and operators
The angular momentum operator in the ladder representation is
Compared to the operator defined in the main text, there is a factor of in the formula here. The harmonic potential in the ladder representation is
The momentum terms in the ladder representation simplify to
We can simplify by noting
Putting the pieces together, we find
Defining the total mode operator we can simplify the Hamiltonian even further:
It is worth pointing out that since and the whole Hamiltonian satisfies Therefore, eigenstates of can be labelled by angular momentum, which we will demonstrate explicitly later.
S2.2
For completeness, we now handle the case. We always assume for otherwise there is no twist. The Hamiltonian now is just As mentioned previously, the operators defined before degenerate and the ladder operators and no longer couple to the coordinates and So this construction does not work for the case. However, a very similar construction does work. Essentially, we define a new operator analogous to with replaced by just and a new operator analogous to also with replaced by just (this replacement is permissible since has units of )
where We inspect the commutation relations: and Using these relations, we find
Now defining we can write the Hamiltonian as
We note that Eq. (S27) is identical to Eq. (S24) with and the appropriate replacements of operators and constants. This is a nice consistency check. However, it is worth pointing out that while Eq. (S24) depends on the sign of Eq. (S27) does not depend on the sign of at all since all the factors appear as This is because when there is no propagation along the vertical direction, so the corresponding eigenstates cannot distinguish between a left-handed twist and a right-handed twist. With that important distinction made, one can now solve Eq. (S27) using the same techniques used for Eq. (S24). This completes the construction of the Landau-level representation of the free-particle Hamiltonian in all relevant scenarios.
S3 Diagonalization of the (Almost) Free-Particle Hamiltonian
S3.1 Special-Function Approach
There are different ways to diagonalize Eq. (S12). Here, we take the most elementary approach using special functions. It is simplest to work in polar coordinates where the Hamiltonian takes the form
where and One can choose to be in either the twisted frame or laboratory frame according to the discussion in Sec. S1.2. The permissible eigenfunctions and energies are
where and are the vertical and lateral sizes of the box regularization employed to normalize the eigenstates, and is a Bessel function of the first kind. It is important to note that these functions are only marginally normalizable (there are other solutions but those are not even marginally normalizable). This makes sense since we are dealing with a free particle without any confining potential. These solutions are known as Bessel beams in the context of electron vortex states [schattschneider2011theory]. The Bessel-function basis is qualitatively very different from the basis states of the operators defined in Sec. S2. While the former is extended, the latter is exponentially well-localized. To see this, we observe that basis states in Sec. S2 are:
where is the reference state annihilated by both and In real space, this reference state is just a Gaussian, and it has zero angular momentum:
To find a general state, let us rewrite the operators in modified complex coordinates
where we have adopted a modified convention for the complex coordinates and Wirtinger derivatives:
We observe the action of these operators on the state:
A technical clarification is in order. The appearance of does not originate directly from the algebra above. We have inserted this by hand to ensure that the resulting state remains an eigenstate of with the same eigenvalue with which we began. This is necessary because the and operators are only defined within a sector. Consequently, it cannot map states outside of that sector. If we insist that the extra exponential factors not be added ad hoc, we can modify the definition of the and operators to explicitly account for them. To do this, we note that adds unit of angular momentum. So, it should become, in the real space representation, On the other hand, depletes unit of angular momentum. So, it should become, in the real space representation, These phases clearly preserve the commutation relations, and do not change the form of the Hamiltonian. The only effect comes when writing down the wavefunctions. These phases ensure that states at with remain inside the sector when acted upon repeatedly by and We shall leave these phases off in the primary discussion, multiplying them in only when necessary. In the twisted frame, when working with and not these phases are never necessary because is completely decoupled from the in-plane degrees of freedom. This technical remark applies equally well to the case of . Now, Eq. (S34) can be simplified by the use of associated Laguerre polynomials
We thus find [robinett2006quantum]
where and The case can be derived analogously. This explicitly shows that the basis states are exponentially localized. Of course, these basis states are not eigenstates of the free-particle Hamiltonian. It seems rather formidable to write eigenstates of the Hamiltonian in this basis as we know they are delocalized Bessel functions. However, the localized basis is convenient, especially for potentially describing crystalline lattices. We therefore would like to make connection between the two. To make progress, we must weakly confine the free electrons. Presumably, in a crystal, there is some confining potential present. So, this addition is physically motivated. In fact, in order to normalize the Bessel functions, we need to impose a large box; therefore, in any case, some type of regularization is needed. The simplest confining potential is an isotropic trap:
where we define a confining length scale For definiteness, let us choose (actually, to simplify the algebra, one can also assume that and individually). Unlike before, we do not need to keep track of the sign of This confining potential, being isotropic, has the same form in either reference frame. The Hamiltonian now is
We can write operators exactly as before
where All the preceding considerations go through effortlessly. In particular, we have and Here, In terms of these new operators, the Hamiltonian is
We can read off the energies immediately
where and are the total mode number and angular momentum defined similar to before. The eigenstates are the same Gauss-Laguerre functions as in Eq. (S36) with replaced by respectively. This approach also makes it apparent that no special treatment is required for the case.
S3.2 Bogoliubov Transformation
A standard approach to solve quadratic bosonic Hamiltonians is using Bogoliubov transformations [colpa1978diagonalization, van1980note, Gurarie2003Bosonic, Complex2005Rossignoli]. Here, we use such a transformation to show that we can get exactly Eq. (S41), confirming the correctness of the result. The quadratic part of the kinetic energy can be rewritten as follows:
We note that this Hamiltonian has a block-diagonal structure because there are no terms of the sort and . A bosonic Bogoliubov transformation is defined by properties of a symplectic group. Williamson’s Theorem states that for any real positive-definite matrix, one can find a symplectic transformation such that the symplectic eigenvalues are positive-definite [nicacio2021williamson]. Unfortunately, the matrix is not positive-definite. Therefore, we need to regularize it by replacing the off-diagonal elements by This regularization turns out to be equivalent to imposing a confining potential, as we will show. We now implement the Bogoliubov transformation [Gurarie2003Bosonic]:
where we assume that and are real numbers. Next, we impose commutation relations on and , which are satisfied if we choose
The last condition can be satisfied by choosing and where is a real parameter. In terms of these new variables, we have
This matrix can be made diagonal by choosing
As long as is real and so are and In this case, the inverse transformations are
So is the Hermitian adjoint of and is the Hermitian adjoint of as desired. The quadratic Hamiltonian is now diagonal
Next, we need the term:
Therefore, in terms of these Bogoliubov operators, we have the following Hamiltonian
We notice that the angular momentum operator now has the same form as before
The eigenstates and energies are given by
Here, is an eigenstate of the number operator and likewise for An alternative way to label states is by their total quanta and angular momentum [footnote1]:
In either case, the eigenstates are labeled by three quantum numbers or because we are in three dimensions. This is true no matter the regularization employed because a Bogoliubov transformation must preserve mode number. Our system is a two-mode bosonic before and after performing a Bogoliubov transformation. The spectrum in Eq. (S53) is almost exactly the spectrum in Eq. (S41). To make them agree exactly, all we have to do is relate to the confining potential. In terms of the original ladder operators, the Hamiltonian with the confining potential becomes
where is defined differently from before (hence the prime). From this, we read off
and the prefactor in the energy becomes exactly as before. To calculate the wavefunctions, we compute
By noting that and we can simplify significantly
This shows that apart from factors of the and operators are the same as the and operators. It is worth pointing out that unlike in the definitions of and where the appearance of is crucial to ensure the commutation relations (because there, and are multiplied by ), the existence of here is a mere gauge choice since the operators are defined primarily with respect to the confining momentum (whose sign is unchanged) and not (whose sign does change). Therefore, and are formally equivalent up to an unimportant gauge choice, confirming that our diagonalization is correct.
S3.3 Energy Spectrum
We end this section by studying the energy spectrum. For convenience, we scale momenta by and energies by (assuming that ). We find the renormalized energy is
To avoid too many variables, we continue to use and to mean both dimensionless and dimensionful quantities. It should be clear from context which definition is employed in a particular instance. Recalling that and where and can be any non-negative integer, we find that must be an integer. This implies that i.e. for a given must be at least as big as Furthermore, and must be of the same parity since the sum and difference have to be divisible by Therefore, where For a fixed the range of is bounded: The bounds are saturated when we have only Bogoliubov quasiparticles of one type. Some representative spectra are shown in Fig. S2. From there, we first observe that if all values of are included, the spectra are dense in energy. This must be the case since we are plotting a four-dimensional data set projected onto a two-dimensional surface. In other words, because the energy is characterized by three quantum numbers, we know that the spectrum is dense when plotted against just one of them (in this case, ). For a fixed the minimum energy For a fixed and the band dispersion is quadratic in with vertex centered at For a fixed the spectrum at different values of are just simply shifted upward in energy by If we are interested in an energy window then roughly only modes with contribute to the physics. It makes sense that the weaker the confinement potential, the more modes must be included in the energy window.
S4 Effect of Saddle-Point Potential Energy
We now consider the effect of a saddle-point potential energy on the energy spectrum of the weakly-confined particles. The potential energy has the following representation in the twisted frame
In the laboratory frame, this same potential energy has a more complicated form
It is straightforward check that in the twisted frame, commutes with since it is independent of In the laboratory frame, commutes with (again, this shows that and are the same operator; so if commutes with it in one frame, it must commute in any other frame). So we can write in each or sector. We can now rewrite in terms of the ladder operators. First, let us do that for the and operators. Recalling that and we find Consequently, we find, in either frame,
In terms of the and operators or the and operators, the potential energy is the same with only one replacement
Because the and operators do not carry cumbersome factors of we choose to write the matrix elements of in terms of these operators. First, we note the usual actions of the operators on the basis states:
Then, writing we have
Evaluating actions on the quadratic terms, we find
We observe by explicit calculation that this potential mixes angular momenta which differ by units. So the spectrum partitions into a mirror-even sector and a mirror-odd sector. Explicitly, the non-zero matrix elements in dimensionless units are
Using these matrix elements, we calculate the energy spectra of particles in the saddle-point potential energy, weakly regularized by the isotropic confining potential. We only include up to a certain maximum value . Physically, this means that there are “hard walls” which prohibit the existence of states beyond a certain radius. In the first calculation, we only include In this case, we find, as shown in Fig. S3, the development of four “edge” states becomes clear as the strength of the saddle-point potential energy increases. If, we include all values up to then the energy spectrum is shown in Fig. S4. There, we also see the “edge” states, but they only become clear for large values of
S5 Boundary Modes of Hofstadter Networks
In the main text, we discussed the bulk spectra of the Hofstadter networks. These networks consist of a periodic array of saddle points described by an intra-saddle-point tunneling amplitude . Tunneling across neighboring saddle points proceeds the inter-saddle-point tunneling amplitude . This is illustrated in Fig. 4 of the main text. In Figs. S5 and S6, we show the spectrum for the bulk and finite-width ribbons of the triangular Hofstadter network, to illustrate the boundary modes appearing in the bulk Hofstadter gaps. We consider both the limit of strong (Fig. S5) and weak (Fig. S6) inter-saddle-point tunneling using the same parameter values as in the main text.